In the right triangle, where a=4 and b=6, find the length of the side not given. what is the value of c ? and then give exact answer to three decimal places.
I appreciate some direction, thank you

Found 2 solutions by Alan3354, MRperkins:
Answer by Alan3354(69443)

In the right triangle, where a=4 and b=6, find the length of the side not given. what is the value of c ? and then give exact answer to three decimal places.
-------------
Assuming c is the hypotenuse, the side opposite the right angle:
c^2 = a^2 + b^2
c^2 = 16 + 36
c^2 = 52
c = sqrt(52) = 2*sqrt(13) exactly
c =~ 7.211

Answer by MRperkins(300)

Use the Pythagorean Theorem: a^2 + b^2 = c^2
If you need help with squares and square roots, then email me at justin.sheppard.tech@hotmail.com and I will be happy to help.
.
Justin